1 litre mixture of CO and CO2 is taken. This is passed through a tube containing red hot charcoal. The 


volume now becomes 1.6 litre. The volumes are measured under the same conditions. Find the 


composition of mixture by volume.


Grade:11

1 Answers

Naveen Kumar
9 years ago
C(s).+ CO2(g)...........>2CO(g)
So all the CO2 being reduced by red hot carbon and make CO gas
let x is the amount of volume of CO2, hence (1-x) litre is the amount of Co present initially.
From equation we can see that one mole of CO2 will produce 2 mole of CO
So,x volume of CO2(g) will give 2x extra CO(g).
So total volume =1-x+2x=1+x=1.6
Hence x=0.6litre=volume of CO2(g) initially, so 60% Co2 and 40% Co(g) initially in the mixture